OpenStudy (sbentleyaz):

find the limit as x approaches infinity of x^20/e^x also written as : \[\lim_{x \rightarrow \infty} x ^{20}/e ^{x}\] a) 0/0 b)1 c)0 d)1/e

OpenStudy (retireed):

e^x will dominate x^20, so you will have large number divided by infinity which is like.... 4/10,000,000 which is essentially zero

OpenStudy (sbentleyaz):

so it would be c which is 0 and not a which is 0/0

OpenStudy (eliesaab):

If you do L'Hospital's Rule 20 times, you will get 20! in the nummerator and e^x in the denominator, so the limit is zero \[ \lim_{x->\infty}\frac {20!}{e^x}=0 \]

OpenStudy (retireed):

It would be c, remember dividing anything by zero, even zero, is undefined.